# Fast and Efficient Matching Algorithm with Deadline Instances

### v2

@@
Zhao Song zsong@adobe.com. Adobe Research.    Weixin Wang weixinw1@uci.edu. UC Irvine.    Chenbo Yin chenboyin1@gmail.com. University of Washington.    Junze Yin junze@bu.edu. Boston University.
The online weighted matching problem is a fundamental problem in machine learning due to its numerous applications. Despite many efforts in this area, existing algorithms are either too slow or don’t take $\mathrm{deadline}$ (the longest time a node can be matched) into account. In this paper, we introduce a market model with $\mathrm{deadline}$ first. Next, we present our two optimized algorithms (FastGreedy and FastPostponedGreedy) and offer theoretical proof of the time complexity and correctness of our algorithms. In FastGreedy algorithm, we have already known if a node is a buyer or a seller. But in FastPostponedGreedy algorithm, the status of each node is unknown at first. Then, we generalize a sketching matrix to run the original and our algorithms on both real data sets and synthetic data sets. Let $\epsilon\in(0,0.1)$ denote the relative error of the real weight of each edge. The competitive ratio of original Greedy and PostponedGreedy is $\frac{1}{2}$ and $\frac{1}{4}$ respectively. Based on these two original algorithms, we proposed FastGreedy and FastPostponedGreedy algorithms and the competitive ratio of them is $\frac{1-\epsilon}{2}$ and $\frac{1-\epsilon}{4}$ respectively. At the same time, our algorithms run faster than the original two algorithms. Given $n$ nodes in $\mathbb{R}^{d}$, we decrease the time complexity from $O(nd)$ to $\widetilde{O}(\epsilon^{-2}\cdot(n+d))$.
@@ 1 Introduction
The online weighted matching problem is a fundamental problem with numerous applications, e.g. matching jobs to new graduates [14], matching customers to commodities [2], matching users to ads [34]. Motivated by these applications, researchers have spent decades designing algorithms to solve this problem [25, 38, 30, 36, 37, 3, 12, 19, 48, 35]. Let $n$ denote the number of items that can be very large in real-world settings. For example, TikTok pushes billions of advertisements every day. Thus, it is necessary to find a method to solve this matching problem quickly. Finding the maximum weight is one of the processes of this method. Our goal is to provide faster algorithms to solve the online weighted matching problem with $\mathrm{deadline}$ by optimizing the part of finding the maximum weight.
In a real-world setting, the weight of the edge between each pair of nodes can show the relationship between two nodes (for example, the higher the weight is, the more likely the buyer wants to buy the product). Each node may have multiple entries to describe its attribute, such as the times of searching for a specific kind of merchandise. And there might exist a time limit called $\mathrm{deadline}$ to confine the longest time a node can be matched. For example, most meat and milk in Amazon need to be sold before it goes bad. This shows the significance of solving the online weighted matching problem with $\mathrm{deadline}$. Let $d$ be the original node dimension and $\mathrm{dl}$ denote $\mathrm{deadline}$. Sometimes the weight is already provided, but we offer a new method to calculate the weight of the edge. This helps accelerate the process of finding the maximum weight.
@@ 1.1 Related Work
Online weighted bipartite matching is a fundamental problem studied by many researchers. [31] provided basic and classical algorithms to solve this problem. Many scientists provided the solution to this problem under various situations [30, 35, 9, 1, 10, 49, 23, 12]. Our work is different from theirs, because our model considered $\mathrm{deadline}$, and our algorithms are $10$ to $20$ times faster than the original algorithms. Moreover, we believe our optimized version is very valuable since it’s very fast and can be extended to other situations with a time limit.
@@
In Section 2, we provide the definition and basic concepts of the model we use. We provide our two optimized algorithms and their proof of correctness in Section 3. We use experiments to justify the advantages and the correctness of our algorithms in Section 4. At last, we draw our conclusion in Section 5.
2 Preliminary
In this section, we first introduce the notations we use in this paper. Then, in Section 2.1, we give the formal definitions related to the bipartite graph. In Section 2.2, we present a useful lemma.
Notations.
We use $\|\cdot\|_{2}$ to denote $\ell_{2}$ norm. For any function $f$, we use $\widetilde{O}(f)$ to denote $f\cdot\operatorname{poly}(\log f)$. For integer $n$, we use $[n]$ to denote $\{1,2,\dots,n\}$. For a set $S$, we use $|S|$ to denote its cardinality.
@@ 2.1 Model
Let $G=(V_{1},V_{2},E)$ denote a bipartite graph with $|V_{1}|=|V_{2}|$. We say $S\subset E$ is a matching if
  • $|S|=|V_{1}|$
  • For each vertex $v\in V_{1}$, there is exactly one edge $e$ in $S$ such that $v$ is one of the vertex in $e$.
  • For each vertex $u\in V_{2}$, there is exactly one edge $e$ in $S$ such that $u$ is one of the vertex in $e$.
Let $w:E\rightarrow\mathbb{R}$ denote a weight function. Let $w_{e}$ denote the weight of edge $e\in E$. We say $w(S)=\sum_{e\in S}w_{e}$ is the weight of matching $S$. Our goal is to make $w(S)$ as large as possible.
@@ 2.2 Useful Lemma
Here, we present a lemma from [26].
The following JL Lemma helps us to create a sketching matrix to accelerate the process of calculating the distance between two points.
@@
For any $X\subset\mathbb{R}^{d}$ of size $n$, there exists an embedding $f:\mathbb{R}^{d}\mapsto\mathbb{R}^{s}$ where $s=O(\epsilon^{-2}\log n)$ such that $(1-\epsilon)\cdot\|x-y\|_{2}\leq\|f(x)-f(y)\|_{2}\leq(1+\epsilon)\cdot\|x-y\|_{2}$.
@@ 3 Algorithm
Proof.
We use $w_{s}$ to denote the current matching value of seller $s$. Let $w_{s,b}$ denote the weight of the edge between the buyer node and $s$-th seller node. Let $i(b)$ be the increment of matching value when buyer $b$ arrives: $i(b)=\max_{s\in{\cal S}}\{w_{s,b},w_{s}\}-w_{s}.$
Let $v_{f}(s)$ be the final matching value for a seller node $s$. If we add all the increments of matching value that every buyer node brings, we will get the sum of every seller node’s matching value. $\sum_{b\in{\cal B}}i(b)=\sum_{s\in{\cal S}}v_{f}(s).$
Let $\mathrm{alg}$ be the weight of the matching in StandardGreedy. Let $\mathrm{alg}=\sum_{s\in{\cal S}}v_{f}(s)$. Let $\mathrm{opt}$ be the optimal matching weight of any matching that can be constructed. Let $\lambda_{i}$ denote the arrival time of node $i$. We need to do the following things to acquire the maximum of $\mathrm{opt}$:
$\mathrm{minimize}$ $~{}\sum_{i\in[n]}a_{i}$
$\mathrm{subject~{}to}$ $~{}a_{i}\geq 0$ $\forall i\in[n]$
$~{}|\lambda_{i}-\lambda_{j}|\leq\mathrm{dl},$ $\forall i<j$
$~{}a_{i}+a_{j}\geq w_{i,j}$ $\forall i,j\in[n].$
We conclude
Lemma 3.2.
Let $w_{i,j}$ denote the real weight of the edge between seller $i$ and buyer $j$, and $\widetilde{w}_{i,j}$ denote the approximated weight. Let $\epsilon\in(0,0.1)$ denote the precision parameter. If there exists a $\alpha$-approximation algorithm for the online weighted matching problem, where $(1-\epsilon)w_{i,j}\leq\widetilde{w}_{i,j}\leq(1+\epsilon)w_{i,j}$ for any seller node $i$ and buyer node $j$, there exists a greedy algorithm with competitive ratio $\alpha(1-\epsilon)$.
Proof.
Since $\epsilon\in(0,0.1)$ and we have $(1-\epsilon)w_{i,j}\leq\widetilde{w}_{i,j}\leq(1+\epsilon)w_{i,j}$. Let $\mathrm{alg}$ and $\mathrm{opt}$ denote the matching weight and optimal weight of the original problem $G=({\cal S},{\cal B},E)$ respectively. Then we define $\mathrm{alg^{\prime}}$ is the matching weight of another problem $G^{\prime}=({\cal S},{\cal B},E^{\prime})$, and our algorithm gives the optimal weight $\mathrm{opt^{\prime}}$.
Since for each edge $\widetilde{w}_{i,j}\leq(1+\epsilon)w_{i,j}$ and $\mathrm{alg}=\sum_{\begin{subarray}{c}i\in{\cal S}\\ j\in{\cal B}\end{subarray}}w_{i,j}$, we can know that $\mathrm{alg^{\prime}}\leq(1+\epsilon)\mathrm{alg}.$ Since for each edge $(1-\epsilon)w_{i,j}\leq\widetilde{w}_{i,j}$ and $\mathrm{opt}=\sum_{\begin{subarray}{c}i\in{\cal S}\\ j\in{\cal B}\end{subarray}}w_{i,j}$, we can know that $\mathrm{opt^{\prime}}\geq(1-\epsilon)\mathrm{opt}.$ Thus,
$\mathrm{alg}\geq$ $~{}\frac{1}{1+\epsilon}\mathrm{alg^{\prime}}\geq~{}\frac{\alpha}{1+\epsilon}\mathrm{opt^{\prime}}$
$\geq$ $~{}\frac{\alpha(1-\epsilon)}{1+\epsilon}\mathrm{opt}\geq~{}\alpha(1-2\epsilon)\mathrm{opt},$
where the first step follows from $\mathrm{alg^{\prime}}\leq(1+\epsilon)\mathrm{alg}$, the second step follows from there exists a $\alpha$-approximation greedy algorithm, the third step follows from $\mathrm{opt^{\prime}}\geq(1-\epsilon)\mathrm{opt}$, and the last step follows from $\frac{\alpha(1-\epsilon)}{1+\epsilon}\geq\frac{\alpha(1-\epsilon)^{2}}{1-\epsilon^{2}}\geq\frac{\alpha(1-\epsilon)^{2}}{2}\geq\alpha(1-2\epsilon)$. Then we can draw our conclusion because we can replace $2\epsilon$ with a new $\epsilon^{\prime}=2\epsilon$ in the last step. ∎
Lemma 3.3.
@@
Proof.
Since the probability of a node being a seller is $1/2$, we can know that
@@
Theorem 3.5 (Informal version of Theorem A.1, Correctness of FastPostponedGreedy in Algorithm 3 and Algorithm 4).
@@
  • Query$(y\in\mathbb{R}^{d})$. Given a query point $y\in\mathbb{R}^{d}$, the Query approximately estimates the Euclidean distances from $y$ to all the data points ${x_{1},x_{2},\dots,x_{n}}\in R^{d}$ in time $O(\epsilon^{-2}(n+d)\log(n/\delta))$. For $\forall i\in[n]$, it provides estimates estimates $\{\widetilde{w}_{i}\}_{i=1}^{n}$ such that:
    $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$
@@
4:     $\widetilde{x}_{1},\widetilde{x}_{2},\dots\widetilde{x}_{n}\in\mathbb{R}^{s}$triangleright Nodes after sketching
5:     $m_{i}$triangleright The vertex matching with vertex $i$
6:     $w_{i}$triangleright Matching value on $x_{i}$
7:     $p$triangleright Matching value
8:     $M$triangleright Sketching matrix
9:     Let $d_{1},d_{2},\dots d_{n}\in\mathcal{N}$ be the deadline for each node $\triangleright$ Each offline point $x_{i}$ can only be matched during time $d_{i}$.
10:     $\text{flag}[n]$triangleright flag[i] decides if node $i$ can be matched.
11:end members
12:procedure Init($x_{1},\dots,x_{n},\epsilon,\delta$)
13:     $p\leftarrow 0$
14:     $s\leftarrow O(\epsilon^{-2}\log(n/\delta))$
15:     for $i=1,2,\dots,s$ do
16:         for $j=1,2,\dots,d$ do
17:              sample $M[i][j]$ from $\{-1/\sqrt{s},+1/\sqrt{s}\}$ each with $1/2$ probability
18:         end for
19:     end for
20:     for $i=1,2,\dots,n$ do
21:         $w_{i}\leftarrow 0$
22:         $\text{flag}[i]\leftarrow 1$
23:         $\widetilde{x}_{i}=Mx_{i}$
24:     end for
25:end procedure
@@
$s=O(\epsilon^{-2}\log(n/\delta))$ is the dimension after transformation. Line 17 is assigning each element of the sketching matrix, and it takes $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$ time since it is a $\mathbb{R}^{s\times d}$ matrix. Line 23 is multiplying the sketching matrix and the vector made up of each coordinate of a node, and it will take $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$. Since we have $n$ nodes to deal with, this whole process will take $O(nsd)=O(\epsilon^{-2}nd\log(n/\delta))$ time.
After carefully analyzing the algorithm, it can be deduced that the overall running time complexity is $O(sd+nsd)=O(\epsilon^{-2}nd\log(n/\delta)).$
@@
Then we will prove the second statement.
We suppose the total matching value $p$ maintains a $\frac{1-\epsilon}{2}$-approximate matching before calling Update. From Lemma 3.9, we can know that after we call Query, we can get $\{\widetilde{w}_{i}\}_{i=1}^{n}$ and for $\forall i\in[n],(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$ with probability $1-\delta$ at least. According to Lemma 3.1 and Lemma 3.2, the competitive ratio of FastGreedy is still $\frac{1-\epsilon}{2}$ after running Query. Therefore, the total matching value $p$ still maintains a $\frac{1-\epsilon}{2}$-approximate matching with probability at least $1-\delta$. ∎
@@
The procedure Query (Algorithm 2) in Theorem 3.6 runs in time $O(\epsilon^{-2}(n+d)\log(n/\delta))$. For $\forall i\in[n]$, it provides estimates estimates $\{\widetilde{w}_{i}\}_{i=1}^{n}$ such that:
$(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$
with probability at least $1-\delta$.
@@
According to Lemma 2.5, we create a sketching matrix $M\in\mathbb{R}^{s\times d}$, let $f(x)=Mx$ and $w_{i}=\|y-x_{i}\|_{2}$ denote the real matching weight of the edge between seller node $i$ and buyer node $y$, and $\widetilde{w}_{i}=\|f(y)-f(x_{i})\|_{2}$ denote the approximated weight of the edge between seller node $i$ and buyer node $y$, then there will be $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$ for $\forall i\in[n]$. Since the failure parameter is $\delta$, for $i\in[n]$ it will provide
$(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$
with probability at least $1-\delta$. ∎
Lemma 3.10 (Informal version of Lemma A.2).
@@
Lemma 3.11 (Informal version of Lemma A.3, Space storage for FastGreedy in Algorithm 1 and Algorithm 2).
@@
4:     $m_{i}$triangleright The vertex matching with vertex $i$
5:     $w_{i}$triangleright Matching value on $x_{i}$
6:     $p$triangleright Matching value
7:     $M$triangleright Sketching matrix
8:     Let $d_{1},d_{2},\dots d_{n}\in\mathcal{N}$ be the deadline for each node $\triangleright$ Each offline point $x_{i}$ can only be matched during time $d_{i}$.
9:end members
10:procedure Init($\epsilon,\delta$)
11:     $p\leftarrow 0$
12:     $s\leftarrow O(\epsilon^{-2}\log(n/\delta))$
13:     ${\cal S}_{0}\leftarrow\emptyset$.
14:     ${\cal B}_{0}\leftarrow\emptyset$.
15:     for $i=1,2,\dots,s$ do
16:         for $j=1,2,\dots,d$ do
17:              sample $M[i][j]$ from $\{-1/\sqrt{s},+1/\sqrt{s}\}$ each with $1/2$ probability
18:         end for
19:     end for
20:end procedure
@@
$s=O(\epsilon^{-2}\log(n/\delta))$ is the dimension after transformation. Line 17 is assigning each element of the sketching matrix, and it takes $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$ time since it is a $\mathbb{R}^{s\times d}$ matrix. After carefully analyzing the algorithm, it can be deduced that the overall running time complexity is $O(sd)=O(\epsilon^{-2}d\log(n/\delta)).$
Lemma 3.14 (Informal version of Lemma A.4).
@@
Lemma 3.15 (Informal version of Lemma A.5).
The procedure Query (Algorithm 4) in Theorem 3.12 runs in time $O(\epsilon^{-2}n\log(n/\delta))$. For $\forall j\in[n]$, it provides estimates $\{\widetilde{w}_{j}\}_{j=1}^{n}$ such that: $(1-\epsilon)\|s_{x_{j}}-b_{x_{i}}\|_{2}\leq\widetilde{w}_{j}\leq(1+\epsilon)\|s_{x_{j}}-b_{x_{i}}\|_{2}$ with probability at least $1-\delta$.
Lemma 3.16 (Informal version of Lemma A.6).
Lemma 3.17 (Informal version of Lemma A.7, Space storage for FastPostponedGreedy in Algorithm 3 and Algorithm 4).

### v3

@@
Zhao Song magic.linuxkde@gmail.com. Simons Institute for the Theory of Computing, UC Berkeley.    Weixin Wang wwang176@jh.edu. Johns Hopkins University.    Chenbo Yin chenboyin1@gmail.com. University of Texas at Austin.    Junze Yin jy158@rice.edu. Rice University.
The online weighted matching problem is a fundamental problem in machine learning due to its numerous applications. Despite many efforts in this area, existing algorithms are either too slow or don’t take $\mathrm{deadline}$ (the longest time a node can be matched) into account. In this paper, we introduce a market model with $\mathrm{deadline}$ first. Next, we present our two optimized algorithms (FastGreedy and FastPostponedGreedy) and offer theoretical proof of the time complexity and correctness of our algorithms. In FastGreedy algorithm, we have already known if a node is a buyer or a seller. But in FastPostponedGreedy algorithm, the status of each node is unknown at first. Then, we generalize a sketching matrix to run the original and our algorithms on both real data sets and synthetic data sets. Let $\epsilon\in(0,0.1)$ denote the relative error of the real weight of each edge. The competitive ratio of original Greedy and PostponedGreedy is $\frac{1}{2}$ and $\frac{1}{4}$ respectively. Based on these two original algorithms, we proposed FastGreedy and FastPostponedGreedy algorithms and the competitive ratio of them is $\frac{1-\epsilon}{2}$ and $\frac{1-\epsilon}{4}$ respectively. At the same time, our algorithms run faster than the original two algorithms. Given $n$ nodes in $\mathbb{R}^{d}$, we decrease the time complexity from $O(nd)$ to $\widetilde{O}(\epsilon^{-2}\cdot(n+d))$, where for any function $f$, we use $\widetilde{O}(f)$ to denote $f\cdot\operatorname{poly}(\log f)$.
@@ 1 Introduction
The online weighted matching problem is a fundamental problem with numerous applications, e.g. matching jobs to new graduates [33], matching customers to commodities [3], matching users to ads [83]. Motivated by these applications, researchers have spent decades designing algorithms to solve this problem [49, 87, 56, 85, 86, 6, 31, 39, 107, 84]. Let $n$ denote the number of items that can be very large in real-world settings. For example, TikTok pushes billions of advertisements every day. Thus, it is necessary to find a method to solve this matching problem quickly. Finding the maximum weight is one of the processes of this method. Our goal is to provide faster algorithms to solve the online weighted matching problem with $\mathrm{deadline}$ by optimizing the part of finding the maximum weight. In a real-world setting, the weight of the edge between each pair of nodes can show the relationship between two nodes (for example, the higher the weight is, the more likely the buyer wants to buy the product). Each node may have multiple entries to describe its attribute, such as the times of searching for a specific kind of merchandise. And there might exist a time limit called $\mathrm{deadline}$ to confine the longest time a node can be matched. For example, most meat and milk in Amazon need to be sold before it goes bad. This shows the significance of solving the online weighted matching problem with $\mathrm{deadline}$. Let $d$ be the original node dimension and $\mathrm{dl}$ denote $\mathrm{deadline}$. Sometimes the weight is already provided, but we offer a new method to calculate the weight of the edge. This helps accelerate the process of finding the maximum weight.
@@ 1.1 Related Work
Online weighted bipartite matching is a fundamental problem that has been studied by many researchers. [57] provided basic and classical algorithms to solve this problem. Our experimental results demonstrate that our optimized version achieves 10-20x speedup while maintaining similar matching quality, making it valuable for time-constrained matching scenarios, like [56, 84, 20, 2, 21, 110, 46, 31].
@@
In Section 2, we introduce the model we use. We provide our two optimized algorithms and their proof of correctness in Section 3. We use experiments to justify the advantages and the correctness of our algorithms in Section 4. At last, we draw our conclusion in Section 5.
2 Preliminaries
Notation.
We use $\|\cdot\|_{2}$ to denote $\ell_{2}$ norm. For any function $f$, we use $\widetilde{O}(f)$ to denote $f\cdot\operatorname{poly}(\log f)$, where $\operatorname{poly}(\log f)$ refers to the polynomial of $\log f$. For integer $n$, we use $[n]$ to denote $\{1,2,\dots,n\}$. For a set $S$, we use $|S|$ to denote its cardinality.
@@ 2.1 Model
Let $G=(V_{1},V_{2},E)$ denote a bipartite graph with $|V_{1}|=|V_{2}|$. We say $S\subset E$ is a matching if $|S|=|V_{1}|$, for each vertex $v\in V_{1}$ there is exactly one edge $e$ in $S$ such that $v$ is one of the vertexes in $e$, and for each vertex $u\in V_{2}$ there is exactly one edge $e$ in $S$ such that $u$ is one of the vertexes in $e$. Let $w:E\rightarrow\mathbb{R}$ denote a weight function. Let $w_{e}$ denote the weight of edge $e\in E$. We say $w(S)=\sum_{e\in S}w_{e}$ is the weight of matching $S$. Our goal is to make $w(S)$ as large as possible.
@@ 2.2 Useful Lemma
JL Lemma creates a sketching matrix to accelerate the process of calculating the distance between two points.
@@
For any $X\subset\mathbb{R}^{d}$ of size $n$, there exists an embedding $f:\mathbb{R}^{d}\to\mathbb{R}^{s}$ where $s=O(\epsilon^{-2}\log n)$ such that $(1-\epsilon)\cdot\|x-y\|_{2}\leq\|f(x)-f(y)\|_{2}\leq(1+\epsilon)\cdot\|x-y\|_{2}$, where $x,y\in X$.
@@ 3 Algorithm
Lemma 3.2 (Restatement of Lemma A.2).
Let $w_{i,j}$ denote the real weight of the edge between seller $i$ and buyer $j$, and $\widetilde{w}_{i,j}$ denote the approximated weight. Let $\epsilon\in(0,0.1)$ denote the precision parameter. If for all $\epsilon$, there exists an $\alpha$-approximation algorithm for the online weighted matching problem and a $\delta>0$, where $(1-\delta)w_{i,j}\leq\widetilde{w}_{i,j}\leq(1+\delta)w_{i,j}$ for any seller node $i$ and buyer node $j$, there exists a greedy algorithm with competitive ratio $\alpha(1-\epsilon)$.
Lemma 3.3 (Restatement of Lemma A.3).
@@
Theorem 3.5 (Restatement of Theorem A.4, Correctness of FastPostponedGreedy in Algorithm 3 and Algorithm 4).
@@
  • Query$(y\in\mathbb{R}^{d})$. Given a query point $y\in\mathbb{R}^{d}$, the Query approximately estimates the Euclidean distances from $y$ to all the data points ${x_{1},x_{2},\dots,x_{n}}\in R^{d}$ in time $O(\epsilon^{-2}(n+d)\log(n/\delta))$. For $\forall i\in[n]$, it provides estimates estimates $\{\widetilde{w}_{i}\}_{i=1}^{n}$ such that: $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$.
@@
3:     $x_{1},x_{2},\dots x_{n}\in\mathbb{R}^{d}$ are nodes in the market, and $\widetilde{x}_{1},\widetilde{x}_{2},\dots\widetilde{x}_{n}\in\mathbb{R}^{s}$ are nodes after sketching.
4:     $m_{i}$ is the vertex matching with vertex $i$, and $w_{i}$ is the matching value on $x_{i}$.
5:     $p$ and $M$ are the matching value and the sketching matrix, respectively.
6:     Let $d_{1},d_{2},\dots d_{n}\in\mathcal{N}$ be the deadline for each node $\triangleright$ Each offline point $x_{i}$ can only be matched during time $d_{i}$.
7:     $\text{flag}[n]$triangleright flag[i] decides if node $i$ can be matched.
8:end members
9:procedure Init($x_{1},\dots,x_{n},\epsilon,\delta$)
10:     $p\leftarrow 0$
11:     $s\leftarrow O(\epsilon^{-2}\log(n/\delta))$
12:     for $i=1,2,\dots,s$ do
13:         for $j=1,2,\dots,d$ do
14:              sample $M[i][j]$ from $\{-1/\sqrt{s},+1/\sqrt{s}\}$ each with $1/2$ probability
15:         end for
16:     end for
17:     for $i=1,2,\dots,n$ do
18:         $w_{i}\leftarrow 0$
19:         $\text{flag}[i]\leftarrow 1$
20:         $\widetilde{x}_{i}=Mx_{i}$
21:     end for
22:end procedure
@@
$s=O(\epsilon^{-2}\log(n/\delta))$ is the dimension after transformation. Line 14 is assigning each element of the sketching matrix, and it takes $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$ time since it is a $\mathbb{R}^{s\times d}$ matrix. Line 20 is multiplying the sketching matrix and the vector made up of each coordinate of a node, and it will take $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$. Since we have $n$ nodes to deal with, this whole process will take $O(nsd)=O(\epsilon^{-2}nd\log(n/\delta))$ time. After carefully analyzing the algorithm, it can be deduced that the overall running time complexity is $O(sd+nsd)=O(\epsilon^{-2}nd\log(n/\delta)).$
@@
Then we will prove the second statement. We suppose the total matching value $p$ maintains a $\frac{1-\epsilon}{2}$-approximate matching before calling Update. From Lemma 3.9, we can know that after we call Query, we can get $\{\widetilde{w}_{i}\}_{i=1}^{n}$ and for $\forall i\in[n],(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$ with probability $1-\delta$ at least. According to Lemma 3.1 and Lemma 3.2, the competitive ratio of FastGreedy is still $\frac{1-\epsilon}{2}$ after running Query. Therefore, the total matching value $p$ still maintains a $\frac{1-\epsilon}{2}$-approximate matching with probability at least $1-\delta$. ∎
@@
The procedure Query (Algorithm 2) in Theorem 3.6 runs in time $O(\epsilon^{-2}(n+d)\log(n/\delta))$. For $\forall i\in[n]$, it provides estimates estimates $\{\widetilde{w}_{i}\}_{i=1}^{n}$ such that: $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$, with probability at least $1-\delta$.
@@
According to Lemma 2.5, we create a sketching matrix $M\in\mathbb{R}^{s\times d}$, let $f(x)=Mx$ and $w_{i}=\|y-x_{i}\|_{2}$ denote the real matching weight of the edge between seller node $i$ and buyer node $y$, and $\widetilde{w}_{i}=\|f(y)-f(x_{i})\|_{2}$ denote the approximated weight of the edge between seller node $i$ and buyer node $y$, then there will be $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$ for $\forall i\in[n]$. Since the failure parameter is $\delta$, for $i\in[n]$ it will provide $(1-\epsilon)\|y-x_{i}\|_{2}\leq\widetilde{w}_{i}\leq(1+\epsilon)\|y-x_{i}\|_{2}$, with probability at least $1-\delta$. ∎
Lemma 3.10 (Restatement of Lemma A.5).
@@
Lemma 3.11 (Restatement of Lemma A.6, Space storage for FastGreedy in Algorithm 1 and Algorithm 2).
@@
4:     $m_{i}$ is the vertex matching with vertex $i$, and $w_{i}$ is the matching value on $x_{i}$.
5:     $p$ and $M$ are the matching value and the sketching matrix, respectively.
6:     Let $d_{1},d_{2},\dots d_{n}\in\mathcal{N}$ be the deadline for each node $\triangleright$ Each offline point $x_{i}$ can only be matched during time $d_{i}$.
7:end members
8:procedure Init($\epsilon,\delta$)
9:     $p\leftarrow 0$
10:     $s\leftarrow O(\epsilon^{-2}\log(n/\delta))$
11:     ${\cal S}_{0}\leftarrow\emptyset$.
12:     ${\cal B}_{0}\leftarrow\emptyset$.
13:     for $i=1,2,\dots,s$ do
14:         for $j=1,2,\dots,d$ do
15:              sample $M[i][j]$ from $\{-1/\sqrt{s},+1/\sqrt{s}\}$ each with $1/2$ probability
16:         end for
17:     end for
18:end procedure
@@
$s=O(\epsilon^{-2}\log(n/\delta))$ is the dimension after transformation. Line 15 is assigning each element of the sketching matrix, and it takes $O(sd)=O(\epsilon^{-2}d\log(n/\delta))$ time since it is a $\mathbb{R}^{s\times d}$ matrix. After carefully analyzing the algorithm, it can be deduced that the overall running time complexity is $O(sd)=O(\epsilon^{-2}d\log(n/\delta)).$
Lemma 3.14 (Restatement of Lemma A.7).
@@
Lemma 3.15 (Restatement of Lemma A.8).
The procedure Query (Algorithm 4) in Theorem 3.12 runs in time $O(\epsilon^{-2}n\log(n/\delta))$. For $\forall j\in[n]$, it provides estimates $\{\widetilde{w}_{j}\}_{j=1}^{n}$ such that: $(1-\epsilon)\|s_{x_{j}}-b_{x_{i}}\|_{2}\leq\widetilde{w}_{j}\leq(1+\epsilon)\|s_{x_{j}}-b_{x_{i}}\|_{2},$ with probability at least $1-\delta$.
Lemma 3.16 (Restatement of Lemma A.9).
The procedure TotalWeight (Algorithm 4) in Theorem 3.12 runs $O(1)$ time. It outputs a $\frac{1-\epsilon}{2}$-approximate matching with probability at least $1-\delta$.
Lemma 3.17 (Restatement of Lemma A.10, Space storage for FastPostponedGreedy in Algorithm 3 and Algorithm 4).